Angle T and Angle S are supplementary. Angle T measures (5x-3)
and Angle S measures (2x+1)
.
Find the measure of each angle.

Answer by ikleyn(52780)

.

Supplementary angles sum up to 180 degrees

       T   +   S    = 180 degrees,

    (5x-3) + (2x+1) = 180

     7x - 2 = 180

     7x = 180 + 2

     7x = 182

      x = 182/7 = 26

      T = 5x-3 = 5*26-3 = 127 degrees,

      S = 2x+1 = 2*26+1 =  53 degrees.

Solved.